OpenStudy (cutiecomittee123):

Derive the trigonometric addition formula for sine sin(a+b)=sinacosb+cosasinb

OpenStudy (princeharryyy):

Use cos(A-B) = cosAcosB + sinAsinB

OpenStudy (princeharryyy):

use 90 deg transformations cos(90-A) = sinA and sin(90-A) = cos(A)

OpenStudy (princeharryyy):

cos(A-B) = cosAcosB + sinAsinB replace A with pi/2-A on both sides cos( pi/2-A -B) = cos (pi/2-A )cosB + sin( pi/2-A )sinB cos( pi/2-(A +B)) = sinAcosB + cosAsinB Sin(A +B) = sinAcosB + cosAsinB

OpenStudy (cutiecomittee123):

Why did you do the part with the 90? How does that have anything to do with the rest of the derivation?

OpenStudy (cutiecomittee123):

@princeharryyy

OpenStudy (princeharryyy):

when the transformations are done >> at 90, sin changes to cos, cos to sin, tan to cot and so on @cutiecomittee123.